OpenStudy (anonymous):

let f(x) = x + 5 and g(x) = x^2. find (g*f)(4) 9 81 20 1

OpenStudy (anonymous):

f(x)=x+5, g(x)=x^2 (g*f)(4)=g{f(4)}=g(4+5)=g(9)=9^2=81

OpenStudy (anonymous):

thank you